he table shows this years nominal GDP and the price level. nstructions: Enter your responses for Nominal and Real GDP rounded to If nominal GDP increases by 2 percent and the price level goes up by 4 percent next year, calculate next years nominal GDP, price level, and real GDP Nominal GDP Price Level Real GDP in Millions) in Millions) This year$ Next year2000 100 $ 1,200.0
Answer #1

This year NGDP = $1200 and it increases by 2% next year. Therefore, next year NGDP = $[1200*(102/100)] = $1224

This year price level(price level measured by the GDP deflator) is 100 and it increases next year by 4% . Therefore, next year price level = 104.

Now, price level = (NGDP/RGDP)*100

Therefore, this year, 104 = (1224/RGDP)*100

Or, (1224/RGDP) = (104/100)

Or, RGDP = (1224*100)/104 = $1176.92

Therefore, next year RGDP = $1176.92